OpenStudy (anonymous):

The point-slope form of the equation of the line that passes through (–5, –1) and (10, –7) is y + 7 = (x – 10). What is the standard form of the equation for this line?

OpenStudy (rowbe98):

The standard form of an equation of a line is Ax + By = C. You will need to convert y+7=(x-10) into that form.

OpenStudy (anonymous):

how do i convert it into that form @rowbe98

OpenStudy (rowbe98):

You need to get all of the variables onto the left side of the equation. Ax and By are simply the x and y variables. Do you know how to do that?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

the equation is y+7=-2/5(x-10) @rowbe98

OpenStudy (anonymous):

that is the actual equation i typed it wrong the first time

OpenStudy (rowbe98):

Okay, no problem! For that equation, you want to use the distributive property on the right side first, to simplify it. So you need to distribute the -2/5 to the (x-10).

OpenStudy (anonymous):

which will make the equation y+7=-2/5x-4

OpenStudy (rowbe98):

Correct! Now you need to get the x on the left side, so all of the variables are on the left.

OpenStudy (anonymous):

y+7+2/5x=-4?

OpenStudy (rowbe98):

Yep! Now move the 7 to the right side, so the variables are isolated. Then write the problem in the proper form, Ax+By=C :)

OpenStudy (anonymous):

so the answer would be y+2/5x=-11 @rowbe98

OpenStudy (rowbe98):

Swap the x and y's positions, since in the standard form x comes first. But yes, you're right! 2/5x+y=-11 :)

OpenStudy (anonymous):

ok thanks

OpenStudy (rowbe98):

You're welcome.